# On Minimax Estimation of Parameters in Softmax-Contaminated Mixture of Experts

### v1

@@ Abstract
footnotetext: Co-first authors, Co-last authors.
@@ 1 Introduction
Mixture of experts (MoE) Jacob_Jordan-1991 ; Jordan-1994 has emerged as a statistical machine learning model that aggregates the power of multiple sub-models. This model consists of two primary components: expert function (or, simply, experts) and a gating network. Experts can be, for example, a feed-forward network (FFN) shazeer2017topk ; dai2024deepseekmoe , a classifier chen2022theory ; nguyen2024general , or a regression model faria2010regression ; kwon_em_2020 . The gating network softly divides the input space into multiple regions where the opinions of some experts are deemed to be more trustworthy than others. This is done by dynamically allocating higher input-dependent weights instead of constant weights to the various experts, making MoE more flexible and adaptive than traditional mixture models Lindsay-1995 . As a consequence, MoE has been leveraged in a wide range of fields, including natural language processing deepseekv3 ; jiang2024mixtral ; grattafiori2024llama3 ; fedus2022switch ; lepikhin_gshard_2021 ; shazeer2017topk , computer vision Riquelme2021scalingvision ; liang_m3vit_2022 , speech recognition You_Speech_MoE ; You_Speech_MoE_2 , multimodal learning han2024fusemoe ; yun2024flexmoe , and reinforcement learning ceron2024rl ; chow_mixture_expert_2023 .
@@
Unlike these applications where all the experts are trainable, parameter-efficient fine-tuning methods such as prefix tuning li-liang-2021-prefix ; le2025revisiting can be interpreted as a mixture of a frozen or pre-trained expert and a trainable prompt expert responsible for learning downstream or more specialized tasks, which we refer to as contaminated MoE throughout this paper. Despite the empirical success of this fine-tuning approach, there is a very limited theoretical understanding of their properties and limitations in the literature. To the best of our knowledge, contaminated MoE has only been previously studied in yan2025contaminated to characterize expert structures achieving the optimal parameter estimation rates. However, the analysis in that work is conducted under a simplified setting where the gating (mixture weight) is independent of the input value, which is a very impractical assumption. To close this gap, in the paper we undertake a through theoretical analysis of the more commonly used softmax-contaminated MoE model, specified in equation (2.1) below, a contaminated MoE model whose gating function takes the form of a soft-maxed linear network.We analyze the issue of identifiability and the convergence properties of the maximum likelihood estimator of the prompt parameters to shed light on the understanding of prompt behavior in prefix tuning methods. A main take-away of our analysis is the potential for the prompt to be exceedingly similar to – and thus to acquire the same knowledge as – the pre-trained model, a situation greatly impacting the estimability of the prompt parameter. To overcome this issue, in Definition 1 we formulate analytical properties of the pre-trained and prompt models, which we refer to as distinguishability, that are guaranteed to rule out excessive overlap between the models and ensure good estimation rates. We make the following contributions.
@@
(i) Distinguishability of the prompt model. In Section 2 we propose a novel notion of distinguishability between the pre-trained and prompt models and illustrate its properties.
@@
(ii) When the distinguishability condition is satisfied, we show in Section 3.1 that the prompt does not converge to the pre-trained model – intuitively, these two models have distinct expertise. In fact, we demonstrate that the convergence rates of the MLE of all the prompt and gating parameters are of parametric order in the sample size $n$, that is, $\widetilde{\mathcal{O}}(n^{-1/2})$. Furthermore, we establish minimax lower bounds on the estimation errors with matching rates, thus showing that the MLE is a minimax rate optimal.
@@
[…]
Table 1: Summary of parameter estimation rates in the softmax-contaminated MoE model. Notice that the rates are in expectation.
@@
Distinguishable $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}(\exp(\tau^{*}))^{-1})$
Non-distinguishable $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}\cdot\|(\Delta\eta^{*},\Delta\nu^{*})\|^{-2})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}\cdot(\|(\Delta\eta^{*},\Delta\nu^{*})\|\exp(\tau^{*}))^{-1})$
@@ 2.1 Problem Setup
Problem setting. Suppose that $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\in\mathcal{X}\times\mathcal{Y}\subset\mathbb{R}^{d}\times\mathbb{R}$ are i.i.d. samples of covariates/response pairs of size $n$. We assume that the input covariates $X_{1},X_{2},\ldots,X_{n}$ are drown in an i.i.d. manner from some known continuous probability distribution on $\mathbb{R}^{d}$ and that the responses are generated according to a softmax-contaminated MoE model, which postulates that the conditional density function of the response given the covariates is given by
@@
Above, the pre-trained model corresponds to as a fixed and known conditional probability density function $f_{0}(\cdot|h_{0}(\cdot,\eta_{0}),\nu_{0})$, parametrized by the pre-trained mean expert function $x\mapsto h_{0}(x,\eta_{0})$ and variance $\nu_{0}$. Meanwhile, the prompt model, denoted as $f(\cdot|h(\cdot,\eta^{*}),\nu^{*})$ is modeled as an unknown Gaussian density function with the prompt mean expert $x\mapsto h(x,\eta^{*})$ and variance $\nu^{*}$. We collect all the unknown parameters of the prompt model into the vector $G_{*}=(\beta^{*},\tau^{*},\eta^{*},\nu^{*})$, belonging to some parameter space $\Xi\subseteq\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{q}\times\mathbb{R}_{>0}$. Note that we allow the values of these parameters to vary with the sample size $n$. However, for notational convenience, we suppress the dependence of $G_{*}$ on $n$ throughout the paper.
Maximum likelihood estimation (MLE). We utilize the maximum likelihood method (Vandegeer-2000, ) to estimate the unknown parameters $G_{*}=(\beta^{*},\tau^{*},\eta^{*},\nu^{*})$ of the softmax-contaminated MoE model (2.1) as follows:
@@
Related work. Mendes et al. mendes2011convergence considered an MoE model where each expert was formulated as a polynomial regression model. Their objective was to address the trade-off between the number of experts and the expert size to obtain the optimal parameter estimation rates. Next, Ho et al. ho2022gaussian took into account the parameter estimation problem for Gaussian MoE models with input-free gating. They demonstrated that when expert functions satisfied an algebraic independence condition, the convergence rates of MLE were optimal of parametric order on the sample size. Conversely, if the expert functions are not algebraic independent, then the parameter estimation rates became inversely proportional to the number of fitted experts. It was not until 2024 that Nguyen et al. nguyen2024deviated investigated a contaminated MoE where a frozen pre-trained model was fine-tuned by a mixture of prompts rather than a single prompt model. However, they imposed two unrealistic assumptions on their model of interest: they equipped the contaminated MoE with input-free gating and kept the ground-truth parameters unchanged with the sample size. Then, Yan et al. yan2025contaminated overcame the second limitation by allowing ground-truth parameters to hinge on the sample size, while the first limitation remained unsolved. Therefore, in this work, our goal is to completely address both limitations by studying the softmax-contaminated MoE in equation (2.1).
@@ 2.2 Fundamental Properties of the Softmax-Contaminated MoE
As mentioned above, when the prompt’s learned skills overlap with those of the pre-trained model, estimating the prompt parameters become challening due to potential non-identifiability. To capture that issue precisely, we introduce an analytic condition called distinguishability in Definition 1.
@@
$\sup_{G_{*}\in\Xi}\mathbb{E}_{p_{G_{*},n}}\Big{[}\mathbb{E}_{X}\left[h\left(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X)\right)\right]\Big{]}\lesssim\sqrt{\log(n)/n}.$ (4)
The above result shows that the density estimator $p_{\widehat{G}_{n}}$ converges to the true density $p_{G_{*}}$ under the Hellinger distance at the near-parametric rate of order $\widetilde{\mathcal{O}}(n^{-1/2})$. To extract from this result a convergence guarantee for the MLE $\widehat{G}_{n}$ itself, we follow a by-now-standard approach in the latest analysis of MoEs; see, e.g., nguyen2023demystifying ; nguyen2024temperature . The main idea is that, if one can exhibit a loss function among parameters, say $D(\widehat{G}_{n},G_{*})$, such that $\mathbb{E}_{p_{G_{*},n}}[D(\widehat{G}_{n},G_{*})]\lesssim\mathbb{E}_{p_{G_{*},n}}\Big{[}\mathbb{E}_{X}\left[h\left(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X)\right)\right]$, then convergence of $\widehat{G}_{n}$ in the expected $D(\cdot,\cdot)$ loss, as well potentially information on the rate of convergence, will follow. See Appendix A for further details. Throughout the rest of the paper, we assume that the tail condition (3) on $f_{0}$ and the distribution of $X$ used in Proposition 3 is in effect.
@@ 3.1 Distinguishable Setting
Suppose that the pre-trained model $f_{0}$ is distinguishable from the prompt model $f$. For almost every $x\in\mathcal{X}$, and for any $\eta\in\mathbb{R}^{q}$, we assume that the Jacobian of the prompt expert function does not vanish, i.e., $\frac{\partial h}{\partial\eta}(x,\eta)\neq 0$. Then, there exists a positive constant $C_{1}$ that depends on $\Xi$ and $f_{0}$ such that the Hellinger lower bound $\mathbb{E}_{X}\left[h(p_{G}\left(\cdot|X),p_{G_{*}}(\cdot|X)\right)\right]\geq C_{1}D_{1}(G,G_{*})$ holds for all parameters $G\in\Xi$. As a result, we obtain
@@
If the pre-trained model $f_{0}$ is distinguishable from the prompt model $f$, then the minimax lower bounds
@@
hold for any $10<r<$ 1, where the infimum is over all estimators taking values in $\Xi$.
@@ 3.2 Non-distinguishable Setting
It should be noted that a key step in obtaining the MLE convergence rates in Theorem 1 is to decompose the density discrepancy $p_{\widehat{G}_{n}}-p_{G_{*}}$ into a combination of linearly independent terms through an appropriate Taylor series expansion of the function $g(y|x;\beta,\eta,\nu):=\exp(\beta^{\top}x)\cdot f(y|h(x,\eta),\nu)$ with respect to its parameters $\beta,\eta,\nu$. This process involves, in particular, higher derivatives of the expert function $h$ with respect to $\eta$, which may not not be algebraically independent. To ensure the linear independence of the terms in the Taylor expansion, we formulate a strong identifiability condition that is indeed sufficient for these purposes.
@@
    $\left\{\frac{\partial h}{\partial\eta^{(u)}},\ \exp(\beta^{\top}x)\frac{\partial h}{\partial\eta^{(u)}},\ x^{(w)}\frac{\partial h}{\partial\eta^{(u)}},\ \frac{\partial^{2}h}{\partial\eta^{(u)}\partial\eta^{(v)}},\ \exp(\beta^{\top}x)\frac{\partial^{2}h}{\partial\eta^{(u)}\partial\eta^{(v)}}\right\}_{u,v\in[q],\ w\in[d]}.$
@@
$-\min\{\exp(\tau),\exp(\tau^{*})\}\left(\|(\Delta\eta,\Delta\nu)\|^{2}\|(\Delta\eta^{*},\Delta\nu^{*})\|^{2}\right)$
@@
Suppose that $f_{0}$ belongs to the family of Gaussian densities and $h_{0}=h$ is strongly identifiable. Then, there exists a positive constant $C_{2}$ that depends on $\Xi,\eta_{0},\nu_{0}$ such that
$\mathbb{E}_{X}\left[h(p_{G}\left(\cdot|X),p_{G_{*}}(\cdot|X)\right)\right]\geq C_{2}D_{2}(G,G_{*})$ holds for all parameters $G$. As a result, we obtain
@@
(ii) Likewise, the convergence rates of the estimators $(\widehat{\beta}_{n},\widehat{\eta}_{n},\widehat{\nu}_{n})$ is also impacted by the convergence rates of the prompt parameters and therefore slower than $\widetilde{\mathcal{O}}(n^{-1/2})$. For example, if $(\Delta\eta^{*},\Delta\nu^{*})$ go to zero at the rate of $\widetilde{\mathcal{O}}(n^{-1/8})$, then the bound (9) indicates that $\widehat{\beta}_{n},\widehat{\eta}_{n},\widehat{\nu}_{n}$ converges to $\beta^{*},\eta^{*},\nu^{*}$ at the rate of $\widetilde{\mathcal{O}}(n^{-3/8})$, respectively. Again, in our numerical experiments below we empirically verify this behavior.
In our final result, whose proof can be found in Appendix A.4, we show that the slower converge rates for the MLE under non-distinguishability from Theorem 8 are in fact essentially minimax optimal.
@@
hold for any sequence $(l_{n})_{n\geq 1}$ and any $10<r<$ 1, , where the infimum is over all estimators $\overline{G}_{n}$ taking values in $\Xi$.

### v2

@@ Abstract
footnotetext: Co-first authors.
@@ 1 Introduction
Mixture of experts (MoE) Jacob_Jordan-1991; Jordan-1994 has emerged as a statistical machine learning model that aggregates the power of multiple sub-models. This model consists of two primary components: expert function (or, simply, expert) and a gating network. Experts can be, for example, a feed-forward network (FFN) shazeer2017topk; dai2024deepseekmoe, a classifier chen2022theory; nguyen2024general, or a regression model faria2010regression; kwon_em_2020. The gating network softly divides the input space into multiple regions where the opinions of some experts are deemed to be more trustworthy than others. This is done by dynamically allocating higher input-dependent weights instead of constant weights to the various experts, making MoE more flexible and adaptive than traditional mixture models Lindsay-1995. As a consequence, MoE has been leveraged in a wide range of fields, including natural language processing deepseekv3; jiang2024mixtral; grattafiori2024llama3; fedus2022switch; lepikhin_gshard_2021; nguyen2025competesmoe; shazeer2017topk, computer vision Riquelme2021scalingvision; liang_m3vit_2022, speech recognition You_Speech_MoE; You_Speech_MoE_2, multimodal learning han2024fusemoe; yun2024flexmoe; nguyen2024hmoe, continual learning le2024mixture; li2025cl, domain adaptation nguyen2025cosine; li2023sparse, and reinforcement learning ceron2024rl; chow_mixture_expert_2023.
@@
Unlike these applications where all experts are trainable, parameter-efficient fine-tuning methods such as prefix tuning li-liang-2021-prefix; le2025revisiting; le2025expressivenessvisualpromptexperts can be interpreted as a mixture of a frozen or pre-trained expert and a trainable prompt expert responsible for learning downstream or more specialized tasks, which we refer to as contaminated MoE throughout this paper. Despite the empirical success of this fine-tuning approach, there is a very limited theoretical understanding of their properties and limitations in the literature. To the best of our knowledge, contaminated MoE has only been previously studied in yan2025contaminated; nguyen2024deviated to characterize expert structures achieving the optimal parameter estimation rates. However, the analysis in that work is conducted under a simplified setting where the gating (mixture weight) is independent of the input value, which is a very impractical assumption. To close this gap, we undertake a thorough theoretical analysis of the more commonly used softmax-contaminated MoE model, specified in equation (2.1) below, a contaminated MoE model whose gating function takes the form of a soft-maxed linear network.We analyze the issue of identifiability and the convergence properties of the maximum likelihood estimator of the prompt parameters to shed light on the understanding of prompt behavior in prefix tuning methods. A main take-away of our analysis is the potential for the prompt to be exceedingly similar to – and thus to acquire the same knowledge as – the pre-trained model, a situation greatly impacting the estimability of the prompt parameter. To overcome this issue, in Definition 1 we formulate analytical properties of the pre-trained and prompt models, which we refer to as distinguishability, that are guaranteed to rule out excessive overlap between the models and ensure good estimation rates. We make the following contributions.
@@
(i) Distinguishability of the prompt model from the pre-trained model. In Section 2, we propose a novel notion of distinguishability between the pre-trained and prompt models and then illustrate its properties.
@@
(ii) When the distinguishability condition is satisfied, we show in Section 3.1 that the prompt does not converge to the pre-trained model – intuitively, these two models have distinct expertise. In fact, we demonstrate that the convergence rates of the MLE of all the prompt and gating parameters are of parametric order in the sample size $n$, that is, $\widetilde{\mathcal{O}}(n^{-1/2})$. Furthermore, we establish minimax lower bounds on the estimation errors with matching rates, thus showing that the convergence rate of MLE is minimax optimal.
@@
Notation. For any $n\in\mathbb{N}$, we let $n]:=\{1,2,\ldots,n\}$. For a vector $u$ we denote with $\|u\|$ its Euclidean norm value. Given any two positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$ and some $C>0$. We further write $a_{n}=\widetilde{\mathcal{O}}(b_{n})$ to denote $a_{n}\lesssim b_{n}\mathrm{polylog}(b_{n})$, where $\mathrm{polylog}(b_{n})$ indicate any term that is polylogarithmic in $b_{n}$. Lastly, for any two densities $p$ and $q$ (dominated by the Lebesgue measure), their squared Hellinger distance is computed as $d_{H}^{2}(p,q):=\frac{1}{2}\int[\sqrt{p(x)}-\sqrt{q(x)}]^{2}dx$, while the total variation distance is given by $d_{V}(p,q):=\frac{1}{2}\int|p(x)-q(x)|dx$.
Table 1: Summary of parameter estimation rates in the softmax-contaminated MoE model. Notice that the rates are in expectation. For the notation, please refer to equations (2.1) and (2). In addition, we also denote $\Delta\eta^{*}:=\eta^{*}-\eta_{0}$ and $\Delta\nu^{*}:=\nu^{*}-\nu_{0}$.
@@
Distinguishable $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}})$
Non-distinguishable $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}\cdot\|(\Delta\eta^{*},\Delta\nu^{*})\|^{-2})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}\cdot\|(\Delta\eta^{*},\Delta\nu^{*})\|^{-1})$
@@ 2.1 Problem Setup
Problem setting. Suppose that $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\in\mathcal{X}\times\mathcal{Y}\subset\mathbb{R}^{d}\times\mathbb{R}$ are i.i.d. samples of covariate-response pairs of size $n$. We assume that the input covariates $X_{1},X_{2},\ldots,X_{n}$ are drown in an i.i.d. manner from some known continuous probability distribution on $\mathbb{R}^{d}$ and that the responses are generated according to a softmax-contaminated MoE model, which postulates that the conditional density function of the response given the covariates is given by
@@
Above, the pre-trained model corresponds to as a fixed and known conditional probability density function $f_{0}(\cdot|h_{0}(\cdot,\eta_{0}),\nu_{0})$, parametrized by the pre-trained mean expert function $x\mapsto h_{0}(x,\eta_{0})$ and variance $\nu_{0}$. Meanwhile, the prompt model, denoted as $f(\cdot|h(\cdot,\eta^{*}),\nu^{*})$ is modeled as an unknown Gaussian density function with the prompt mean expert $x\mapsto h(x,\eta^{*})$ and variance $\nu^{*}$. We collect all the unknown parameters of the prompt model into the vector $G_{*}=(\beta^{*},\tau^{*},\eta^{*},\nu^{*})$, belonging to some parameter space $\Xi\subseteq\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{q}\times\mathbb{R}_{+}$. Note that we allow the values of these parameters to vary with the sample size $n$. However, for notational convenience, we suppress the dependence of $G_{*}$ on $n$ throughout the paper. In addition, it should also be noted that the “probabilistic" MoE model (2.1) can be related to “deterministic" MoE models used in deep learning shazeer2017topk by taking the expectation of the response given the covariate, that is,
$\mathbb{E}[Y|X]=\frac{1}{1+\exp((\beta^{*})^{\top}x+\tau^{*})}\cdot h_{0}(x,\eta_{0})+\frac{\exp((\beta^{*})^{\top}x+\tau^{*})}{1+\exp((\beta^{*})^{\top}x+\tau^{*})}\cdot h(x,\eta^{*}).$
Maximum likelihood estimation (MLE). We utilize the maximum likelihood method (Vandegeer-2000) to estimate the unknown parameters $G_{*}=(\beta^{*},\tau^{*},\eta^{*},\nu^{*})$ of the softmax-contaminated MoE model (2.1) as follows:
@@
Related work. Mendes et al. mendes2011convergence considered an MoE model where each expert was formulated as a polynomial regression model. Their objective was to address the trade-off between the number of experts and the expert size to obtain the optimal parameter estimation rates. Next, Ho et al. ho2022gaussian took into account the parameter estimation problem for Gaussian MoE models with input-free gating. They demonstrated that when expert functions satisfied an algebraic independence condition, the convergence rates of MLE were optimal of parametric order on the sample size. Conversely, if the expert functions are not algebraic independent, then the parameter estimation rates became inversely proportional to the number of fitted experts. These results were then extended to more practical settings of input-dependent gatings, including softmax gating nguyen2023demystifying and sigmoid gating nguyen2024sigmoid, revealing that the latter was more sample-efficient than the former in terms of expert estimation.
It was not until 2024 that Nguyen et al. nguyen2024deviated investigated a contaminated MoE where a frozen pre-trained model was fine-tuned by a mixture of prompts rather than a single prompt model. However, they imposed two unrealistic assumptions on their model of interest: they equipped the contaminated MoE with input-free gating and kept the ground-truth parameters unchanged with the sample size. Then, Yan et al. yan2025contaminated overcame the second limitation by allowing ground-truth parameters to hinge on the sample size as in the case of traditional mixture models do2023deviated, while the first limitation remained unsolved. Therefore, in this work, our goal is to completely address both limitations by studying the softmax-contaminated MoE in equation (2.1).
Challenges. There are three fundamental challenges of our analysis compared to previous work.
1. Uniform convergence rates. We allow ground-truth parameters $G_{*}$ to change with sample size $n$, which is challenging yet closer to practice than the settings in previous works on MoE nguyen2023demystifying; nguyen2024sigmoid, where $G_{*}$ does not change with $n$. Thus, the convergence rates of parameter estimations in our work are uniform rather than point-wise as in those works.
2. Minimax lower bounds. We determine minimax lower bounds under both distinguishable and non-distinguishable settings. Based on these lower bounds, we can claim that our derived convergence rates are optimal. However, no minimax lower bounds are provided in nguyen2023demystifying; nguyen2024sigmoid.
3. Input-dependent gating. The latest work on understanding the contaminated MoE model is yan2025contaminated, but it considers input-free gating in the analysis. On the other hand, in this paper, we take into account softmax gating, which hinges upon the input value. This input-dependence yields several challenges on the convergence of density estimation and parameter estimation.
@@ 2.2 Fundamental Properties of the Softmax-Contaminated MoE
As mentioned above, when the prompt’s learned skills overlap with those of the pre-trained model, estimating the prompt parameters becomes challenging due to potential non-identifiability. To capture that issue accurately, we introduce an analytic condition called distinguishability in Definition 1.
@@
$\sup_{G_{*}\in\Xi}\mathbb{E}_{p_{G_{*},n}}\Big[\mathbb{E}_{X}\left[d_{H}\left(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X)\right)\right]\Big]\lesssim\sqrt{\log(n)/n}.$ (4)
The above result shows that the density estimator $p_{\widehat{G}_{n}}$ converges to the true density $p_{G_{*}}$ under the Hellinger distance at the near-parametric rate of order $\widetilde{\mathcal{O}}(n^{-1/2})$. To extract from this result a convergence guarantee for the MLE $\widehat{G}_{n}$ itself, we follow a by-now-standard approach in the latest analysis of MoEs; see, e.g., nguyen2023demystifying. The main idea is that, if one can exhibit a loss function among parameters, say $D(\widehat{G}_{n},G_{*})$, such that $\mathbb{E}_{p_{G_{*},n}}[D(\widehat{G}_{n},G_{*})]\lesssim\mathbb{E}_{p_{G_{*},n}}\Big[\mathbb{E}_{X}\left[d_{H}\left(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X)\right)\right]$, then convergence of $\widehat{G}_{n}$ in the expected $D(\cdot,\cdot)$ loss, as well potentially information on the rate of convergence, will follow. See Appendix A for further details. Throughout the rest of the paper, we assume that the tail condition (3) on $f_{0}$ and the distribution of $X$ used in Proposition 3 is in effect.
@@ 3.1 Distinguishable Setting
Suppose that the pre-trained model $f_{0}$ is distinguishable from the prompt model $f$. For almost every $x\in\mathcal{X}$, and for any $\eta\in\mathbb{R}^{q}$, we assume that the Jacobian of the prompt expert function does not vanish, i.e., $\frac{\partial h}{\partial\eta}(x,\eta)\neq 0$. Then, there exists a positive constant $C_{1}$ that depends on $\Xi$ and $f_{0}$ such that the Hellinger lower bound $\mathbb{E}_{X}\left[d_{H}(p_{G}\left(\cdot|X),p_{G_{*}}(\cdot|X)\right)\right]\geq C_{1}D_{1}(G,G_{*})$ holds for all parameters $G\in\Xi$. As a result, we obtain
@@
If the pre-trained model $f_{0}$ is distinguishable from the prompt model $f$, then the following minimax lower bounds hold for any $0<r<1$:
@@
where the infimum is over all estimators $\overline{G}_{n}:=(\overline{\beta}_{n},\overline{\tau}_{n},\overline{\eta}_{n},\overline{\nu}_{n})$ taking values in $\Xi$.
@@ 3.2 Non-distinguishable Setting
It should be noted that a key step in obtaining the MLE convergence rates in Theorem 1 is to decompose the density discrepancy $p_{\widehat{G}_{n}}-p_{G_{*}}$ into a combination of linearly independent terms through an appropriate Taylor series expansion of the function $g(y|x;\beta,\eta,\nu):=\exp(\beta^{\top}x)\cdot f(y|h(x,\eta),\nu)$ with respect to its parameters $\beta,\eta,\nu$. This process involves, in particular, higher derivatives of the expert function $h$ with respect to $\eta$, which may not be algebraically independent. To ensure the linear independence of the terms in the Taylor expansion, we formulate a strong identifiability condition that is indeed sufficient for these purposes.
@@
    $\hskip-20.00003pt\left\{\frac{\partial h}{\partial\eta^{(u)}},\ \exp(\beta^{\top}x)\frac{\partial h}{\partial\eta^{(u)}},\ x^{(w)}\frac{\partial h}{\partial\eta^{(u)}},\ \frac{\partial^{2}h}{\partial\eta^{(u)}\partial\eta^{(v)}},\ \exp(\beta^{\top}x)\frac{\partial^{2}h}{\partial\eta^{(u)}\partial\eta^{(v)}}\right\}_{u,v\in[q],\ w\in[d]}.$
@@
$-\min\{\exp(\tau),\exp(\tau^{*})\}\left(\|(\Delta\eta,\Delta\nu)\|^{2}+\|(\Delta\eta^{*},\Delta\nu^{*})\|^{2}\right)$
@@
Suppose that $f_{0}$ belongs to the family of Gaussian densities and $h_{0}=h$. Then, there exists a positive constant $C_{2}$ that depends on $\Xi,\eta_{0},\nu_{0}$ such that $\mathbb{E}_{X}\left[d_{H}(p_{G}\left(\cdot|X),p_{G_{*}}(\cdot|X)\right)\right]\geq C_{2}D_{2}(G,G_{*})$ holds for all parameters $G$. As a result, we obtain
@@
(ii) Likewise, the convergence rates of the estimators $(\widehat{\beta}_{n},\widehat{\eta}_{n},\widehat{\nu}_{n})$ are also impacted by the convergence rates of the prompt parameters and therefore slower than $\widetilde{\mathcal{O}}(n^{-1/2})$. For example, if $(\Delta\eta^{*},\Delta\nu^{*})$ go to zero at the rate of $\widetilde{\mathcal{O}}(n^{-1/8})$, then the bound (9) indicates that $\widehat{\beta}_{n},\widehat{\eta}_{n},\widehat{\nu}_{n}$ converges to $\beta^{*},\eta^{*},\nu^{*}$ at the rate of $\widetilde{\mathcal{O}}(n^{-3/8})$, respectively. Again, in our numerical experiments below we empirically verify this behavior.
In our final result, whose proof can be found in Appendix A.4, we show that the slower converge rates for the MLE under non-distinguishability are in fact essentially minimax optimal.
@@
hold for any sequence $(l_{n})_{n\geq 1}$ and any $0<r<1$, , where the infimum is over all estimators $\overline{G}_{n}$ taking values in $\Xi$.
3.3 Practical Implications
There are two important practical implications for the design of a contaminated MoE model from our theoretical results.
1. Softmax gating is more sample-efficient than input-free gating. We observe that softmax gating yields faster convergence rates of prompt parameter estimation in contaminated MoE than input-free gating in yan2025contaminated. In particular, when using input-free gating, Table 2 reveals that the rates for estimating expert parameters and variance depend on the convergence rate of the gating parameter to zero. By contrast, when using softmax gating, estimation rates for expert parameters and variance become significantly faster as the previous rate dependence disappears. Therefore, our theories encourage the use of softmax gating over input-free gating when tuning contaminated-MoE-based models.
2. Prompt models should have different expertise from pre-trained models. It can be seen from Table 2 that when the prompt model acquires overlapping knowledge with the pre-trained model (non-distinguishable setting), the convergence rates of parameter estimation are slower than when these models have distinct knowledge (distinguishable setting). Thus, our theories advocate using prompt models with different expertise from the pre-trained model.
Table 2: Comparison of parameter estimation rates in input-free-contaminated MoE yan2025contaminated and softmax-contaminated MoE (Ours). Below, we consider gating parameters $\exp(\beta^{*}_{0})$, expert parameters $\eta^{*}$, and variance $\nu^{*}$. In addition, $\lambda^{*}$ denotes the constant weight in input-free-contaminated MoE.
Distinguishable Setting
Gating parameters Expert parameters and Variance
Input-free gating yan2025contaminated $\widetilde{\mathcal{O}}(n^{-1/2})$ $\widetilde{\mathcal{O}}(n^{-1/2}(\lambda^{*})^{-1})$
Softmax gating (Ours) $\widetilde{\mathcal{O}}(n^{-1/2})$ $\widetilde{\mathcal{O}}(n^{-1/2})$
Non-distinguishable Setting
Gating parameters Expert parameters and Variance
Input-free gating yan2025contaminated $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}\cdot\|(\Delta\eta^{*},\Delta\nu^{*})\|^{-2})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}\cdot\|(\Delta\eta^{*},\Delta\nu^{*})\|^{-1}(\lambda^{*})^{-1})$
Softmax gating (Ours) $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}\cdot\|(\Delta\eta^{*},\Delta\nu^{*})\|^{-2})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{2}}\cdot\|(\Delta\eta^{*},\Delta\nu^{*})\|^{-1})$